OpenStudy (anonymous):

Find the y-intercept of the graph of this equation: 4y + 8x = 24 A. (0, 8) B. (0, 6) C. (8, 0) D. (6, 0)

OpenStudy (ivettef365):

to find the y-int, just substitute the x = 0 and what do you get ??

OpenStudy (anonymous):

we will need to put the equation in slope intercept form: y=mx+b and 'b' will be your y intercept so 4y +8x=24 becomes 4y = -8x+24 divide by 4 y = -2x+6

OpenStudy (anonymous):

so its b?

OpenStudy (ivettef365):

yes, b is correct